# CDI3D: Cross-guided Dense-view Interpolation for 3D Reconstruction

### v1

@@ I Introduction
Recent mainstream deep learning-based 3D representations, such as voxels [5, 6, 7], point clouds [8, 9], and implicit functions [10, 11], have been widely studied and applied in various 3D reconstruction tasks. Although these paradigms have demonstrated distinct advantages across various applications, their fundamental limitations persist, such as high memory consumption for voxels, lack of connectivity in point clouds, and complex optimization requirements for implicit functions. In contrast, meshes have gained increasing attention due to their ability to explicitly encode both geometry and topology in a compact and structured form, making them particularly suitable for detailed surface reconstruction and widely used in practical applications. Previous works have explored the generation of meshes directly from images using end-to-end encoder-decoder architectures, where an encoder extracts image features and a decoder predicts the reconstructed mesh [12, 13, 14]. However, such methods face challenges, including high computational demands and difficulties in accurately modeling complex mesh geometry, often leading to suboptimal surface quality and limited generalization.
@@
Meanwhile, advancements in 2D diffusion models [17, 18, 19] and Large Reconstruction Models (LRMs) [20, 21, 22, 23, 24] have opened new avenues for 3D content creation. Several works, including [25, 26, 27, 28, 29, 30, 31], leverage 2D diffusion models to generate 3D content through a Score Distillation Sampling (SDS) pipeline, which optimizes a 3D representation by aligning its rendered 2D projections with the distribution learned by a pre-trained 2D diffusion model. An alternative paradigm involves generating multi-view images using 2D diffusion models, followed by applying reconstruction algorithms to infer 3D geometry and topology from these images [32, 33, 34, 35, 36, 37], resulting in more detailed geometry and texture reconstruction with improved performance. However, current state-of-the-art (SoTA) methods face notable limitations. Most approaches generate a limited number of multi-view images (typically four or six), which constrains the richness of geometric and textural details that can be captured. To address this, recent works such as [38, 15, 16] have proposed video diffusion strategies to directly expand the number of generated views. While these methods mark a step forward, they are often hindered by significant challenges, particularly multi-view inconsistency, where the generated views fail to align cohesively, leading to artifacts and degraded reconstruction quality, as illustrated in Fig. 1 (SV3D [15] and V3D [16]). Furthermore, these approaches come with substantial computational overhead, requiring intensive GPU memory and extended training times, which severely limit their scalability and practical applicability in real-world scenarios.
@@ II Related Works
3D generation from a single image is a long-standing and challenging problem in computer vision. Mainstream learning-based 3D representations include voxels [5, 6, 7], point clouds [8, 9], and implicit functions [10, 11]. While these representations have shown promise, they often struggle to balance efficiency, accuracy, and scalability. Recently, meshes have emerged as a compelling alternative due to their structured and compact nature, which allows for explicit encoding of both geometry and topology. This makes them particularly advantageous for tasks requiring detailed and high-quality surface reconstruction.
Building on the advancements in 3D representations, the remarkable success of diffusion models in 2D image generation [17, 18, 40, 19] has inspired numerous works to explore their potential for 3D generation. By leveraging the powerful generative capabilities of diffusion models, researchers have begun to adapt these techniques to tackle the challenges of 3D content creation. A mainstream approach is directly training 3D generators using 3D ground truth [41, 42, 43, 44, 45]. For instance, [41] and [42] trained diffusion models to directly generate 3D voxels. In [43] and [45], a 3D-aware tri-plane diffusion model is introduced to produce NeRF [46] representations. Nonetheless, 3D diffusion methods tend to be time-consuming during optimization, and often show low quality in terms of texture and geometry. To deal with this, some studies have explored the utilization of 2D diffusion-based generators for 3D generation. DreamFusion [25] was the first to use 2D diffusion models to generate 3D content through a SDS pipeline. Building upon this work, [26, 27, 28, 29, 30, 31] have adopted the SDS pipeline to optimize various 3D representations such as NeRF, mesh, and gaussian splatting [47]. However, performing 3D generation tasks with 2D diffusion models often encounters issues related to multi-view inconsistency, indicating room for improvement.
@@ II-C Large Reconstruction Models
The advent of large-scale 3D datasets [3, 4] has significantly advanced the field of image-to-3D generation, bringing generalized reconstruction models to new heights. LRM [20] was a pioneer that demonstrates the superiority of Transformer [48] backbone in mapping image tokens to predict tri-plane NeRF under multi-view supervision. Building upon this foundation, Instant3D [21] extends the input to multi-view images, largely enhancing the quality of image-to-3D generation through multi-view diffusion models. Inspired by Instant3D, subsequent methods such as LGM [22] and GRM [49] further refine it by replacing NeRF representations with 3D Gassian Splatting [47] to improve the rendering efficiency. Recently, CRM [23] and InstantMesh [24] take advantage of FlexiCubes [50] to improve both efficiency and quality of image-to-3D generation.

### v2

@@ I Introduction
Recent mainstream deep learning-based 3D representations, such as voxels [5, 6], point clouds [7, 8], and implicit functions [9, 10], have been widely studied and applied in various 3D reconstruction tasks. Although these paradigms have demonstrated distinct advantages across various applications, their fundamental limitations persist, such as high memory consumption for voxels, lack of connectivity in point clouds, and complex optimization requirements for implicit functions. In contrast, meshes have gained increasing attention due to their ability to explicitly encode both geometry and topology in a compact and structured form, making them particularly suitable for detailed surface reconstruction and widely used in practical applications. Previous works have explored the generation of meshes directly from images using end-to-end encoder-decoder architectures, where an encoder extracts image features and a decoder predicts the reconstructed mesh [11, 12]. However, such methods face challenges, including high computational demands and difficulties in accurately modeling complex mesh geometry, often leading to suboptimal surface quality and limited generalization.
@@
Meanwhile, advancements in 2D diffusion models [15, 16] and Large Reconstruction Models (LRMs) [17, 18, 19, 20, 21] have opened new avenues for 3D content creation. Several works, including [22, 23, 24, 25, 26, 27], leverage 2D diffusion models to generate 3D content through a Score Distillation Sampling (SDS) pipeline, which optimizes a 3D representation by aligning its rendered 2D projections with the distribution learned by a pre-trained 2D diffusion model. An alternative paradigm involves generating multi-view images using 2D diffusion models, followed by applying reconstruction algorithms to infer 3D geometry and topology from these images [28, 29, 30, 31, 32, 33], resulting in more detailed geometry and texture reconstruction with improved performance. However, current state-of-the-art (SoTA) methods face notable limitations. Most approaches generate a limited number of multi-view images (typically four or six), which constrains the richness of geometric and textural details that can be captured. To address this, recent works such as [34, 13, 14] have proposed video diffusion strategies to directly expand the number of generated views. While these methods mark a step forward, they are often hindered by significant challenges, particularly multi-view inconsistency, where the generated views fail to align cohesively, leading to artifacts and degraded reconstruction quality, as illustrated in Fig. 1 (SV3D [13] and V3D [14]). Furthermore, these approaches come with substantial computational overhead, requiring intensive GPU memory and extended training times, which severely limit their scalability and practical applicability in real-world scenarios.
@@ II Related Works
3D generation from a single image is a long-standing and challenging problem in computer vision. Mainstream learning-based 3D representations include voxels [5, 6], point clouds [7, 8], and implicit functions [9, 10]. While these representations have shown promise, they often struggle to balance efficiency, accuracy, and scalability. Recently, meshes have emerged as a compelling alternative due to their structured and compact nature, which allows for explicit encoding of both geometry and topology. This makes them particularly advantageous for tasks requiring detailed and high-quality surface reconstruction.
Building on the advancements in 3D representations, the remarkable success of diffusion models in 2D image generation [15, 16] has inspired numerous works to explore their potential for 3D generation. By leveraging the powerful generative capabilities of diffusion models, researchers have begun to adapt these techniques to tackle the challenges of 3D content creation. A mainstream approach is directly training 3D generators using 3D ground truth [36, 37, 38, 39, 40]. For instance, [36] and [37] trained diffusion models to directly generate 3D voxels. In [38] and [40], a 3D-aware tri-plane diffusion model is introduced to produce NeRF [41] representations. Nonetheless, 3D diffusion methods tend to be time-consuming during optimization, and often show low quality in terms of texture and geometry. To deal with this, some studies have explored the utilization of 2D diffusion-based generators for 3D generation. DreamFusion [22] was the first to use 2D diffusion models to generate 3D content through a SDS pipeline. Building upon this work, [23, 24, 25, 26, 27] have adopted the SDS pipeline to optimize various 3D representations such as NeRF, mesh, and gaussian splatting [42]. However, performing 3D generation tasks with 2D diffusion models often encounters issues related to multi-view inconsistency, indicating room for improvement.
@@ II-C Large Reconstruction Models
The advent of large-scale 3D datasets [3, 4] has significantly advanced the field of image-to-3D generation, bringing generalized reconstruction models to new heights. LRM [17] was a pioneer that demonstrates the superiority of Transformer [43] backbone in mapping image tokens to predict tri-plane NeRF under multi-view supervision. Building upon this foundation, Instant3D [18] extends the input to multi-view images, largely enhancing the quality of image-to-3D generation through multi-view diffusion models. Inspired by Instant3D, subsequent methods such as LGM [19] and GRM [44] further refine it by replacing NeRF representations with 3D Gassian Splatting [42] to improve the rendering efficiency. Recently, CRM [20] an InstantMesh [21] takes advantage of FlexiCubes [45] to improve both efficiency and quality of image-to-3D generation.